### v3

@@ 5 Department of Astronomy, Westlake University, Hangzhou, Zhejiang 310030, China
A Vertically Orientated Dark Matter Halo Marks a Flip of the Galactic Disk
@@
[…]
@@ 1 Introduction
Determining the exact shape of the DM halo from observations, even for the Milky Way, is a challenge. One major method to probe the DM distribution is to model the motions of tracers, which reflect the underlying gravitational potential. Stellar streams, viewed as simple orbital pathways that trace the halo, played a significant role in constraining the shape of the Milky Way DM halo (Bovy2016ApJ...833...31B; Bowden2015MNRAS.449.1391B; Koposov2010ApJ...712..260K; Palau2023MNRAS.524.2124P; Law2010ApJ...714..229L). However, the most commonly used dynamically cold streams, e.g., Pal 5 and GD–1, trace only the inner halo at $r\lesssim 20$ kpc. And a single stream is easily disturbed (Nibauer2024ApJ...969...55N) and may lead to biassed results (Vasiliev2021MNRAS.501.2279V).
@@
Halo stars - with the full 6D phase-space information observed - span the entire space and, statistically, should exhibit strong power in constraining the underlying 3D DM distribution. However, these stars exist with complex orbital dynamics, necessitating accurate dynamical modelling. At present, the Milky Way halo is analysed using two primary types of dynamical models: the Jeans model and the distribution function (DF)-based dynamical model. These models yield divergent results regarding the shape of the DM halo, ranging from oblate with $q\sim 0.4-0.7$ (Loebman2014ApJ...794..151L; Hattori2021MNRAS.508.5468H; Li2022MNRAS.510.4706L), to nearly spherical with $q\sim 0.9-1$ (Wegg2019MNRAS.485.3296W; Zhang2025), and to prolate with $q\sim 1.3$ (Posti2019A&A...621A..56P). All these models are restricted by axisymmetric assumptions and incorporate ad-hoc assumptions on the intrinsic dynamical structures. Despite extensive research over the years, the true shape of the Milky Way DM halo remains debated (Hunt2025NewAR.10001721H).
@@
Due to the large uncertainty in the DM halo shape of the Milky Way from observations, its link to the spatial arrangement of satellite galaxies has not been seriously considered. The satellite galaxies in the Milky Way do not distribute isotropically but rather form a thin plane vertical to the Galactic disk (Lynden-Bell1976MNRAS.174..695L; Kroupa2005A&A...431..517K; Pawlowski2018MPLA...3330004P; Sawala2023NatAs...7..481S). The thickness of this plane of satellites has been extensively studied; it was previously thought to be rare but has now become more frequently observed through cosmological simulations (Sawala2023NatAs...7..481S). The cause of its perpendicular alignment to the Galactic disk remains uncertain (Vasiliev2021MNRAS.501.2279V). Typically, both the stellar disk and satellite system align parallel to the DM halo, sharing a similar net angular momentum direction, leading to a co-alignment of the stellar disk and satellite plane (Shao2016MNRAS.460.3772S). However, in systems resembling the Milky Way, where the satellite system lies perpendicular to the stellar disk, the DM halo might exhibit a twist, being co-aligned with the stellar disk in the inner regions but aligned vertically in the outer regions (Shao2021MNRAS.504.6033S). Precisely measuring the DM shape as a function of radius could provide essential insights into the origin of the vertical satellite plane and elucidate its relationship with the large-scale matter distribution. Additionally, this suggests that the DM halo of the Milky Way may be triaxial, with a shape that changes from the inner to outer regions, necessitating dynamic models with more flexible DM halos.
@@
Recently, a triaxial Schwarzwald model that numerically represents the DF of tracers by superposition of stellar orbits has been applied to the Milky Way halo, indicating a triaxial, nearly prolate halo with a large tilt angle (Dillamore2025arXiv251000095D). Nevertheless, there exists a strong degeneracy between the axis ratios of the DM halo and the tilt angle. This model was mainly constrained by the shape of a tilted stellar halo (Han2022AJ....164..249H) associated with the Gaia-Sausage-Enceladus (GSE) (Helmi2018Natur.563...85H; Belokurov2018MNRAS.478..611B).
@@
In a previous paper (Zhu2025A&A...703A..43Z), we presented and validated an empirical triaxial orbit-superposition model tailored for the Milky Way halo observed by the Large Sky Area Multi-Object Fibre Spectroscopic Telescope (LAMOST) and the Global Astrometric Interferometer for Astrophysics (Gaia). This approach accommodates a highly flexible triaxial DM halo, where the axis ratios are allowed to vary with radius, and it uses the full 6D phase-space information from the observations. Compared to traditional made-to-measure and Schwarzschild methods, this method did not theoretically sample the orbit libraries or determine the orbit weights by fitting the model to data; instead, the orbit library and orbit weights are entirely determined from the observations. The results are thus highly data-driven.
@@
In this paper, we apply the empirical triaxial orbit-superposition model to the Milky Way halo, utilising 6D phase-space data from LAMOST and Gaia, and robustly determine the 3D DM distribution of the Milky Way, extending to approximately 50 kpc. We then compare the findings with Milky Way analogues from the Auriga and TNG50 cosmological simulations and establish a coherent scenario linking the DM halo shape/orientation and satellite arrangement. The paper is organised in the following way: we describe the details of observational data in Section 2, introduce model construction in Section 3, and show the results in Section 4. We compare the 3D shape of the DM halo and the arrangement of the Milky Way’s satellite system with cosmological simulations in Section 5. We discuss in Section 6 and conclude in Section 7.
@@ 2.1 Data preparation
Figure 1: The density of stars in the parameter phase of circularity $\lambda_{z}$ versus metallicity [Fe/H]. We first remove the disk stars with circularity $\lambda_{z}>0.8$, then also remove the group of stars with $\lambda_{z}>0.3$ and $\rm Fe/H]>-1.2$. The rest is kept as halo stars in our sample.
Figure 2: The K giants in our final sample of smooth halo in the northern hemisphere, colored by the weights of particles obtained by correction of selection function. The sample can be taken as complete in the $R_{\rm gc}-z_{\rm gc}$ space after correction of selection function.
We obtain the 3D positions and 3D velocities of halo stars by combining observations from LAMOST and Gaia. LAMOST has gathered spectroscopic data for approximately 10 million stars within the Milky Way. We used k giants from LAMOST DR9; the distance is accurately measured with a relative uncertainty of $\sim 15\%$ (Zhang2023). By cross-matching with Gaia DR3, we have a sample of 619,284 k-giants with 3D velocity measurements; the typical errors in radial velocity $v_{r}$, tangential velocity $v_{\phi}$, and $v_{\theta}$ are 10, 30, and 30 km/s, respectively. We exclude disk stars chemodynamically to avoid biassing the spatial distribution of the halo. We first remove all the stars on cold disk with $\lambda_{z}>0.8$, then remove the group of stars with $\lambda_{z}>0.3$ and $\rm Fe/H]>-1.2$ which are mostly within 20 kpc, not in dynamical equilibrium and likely to be induced from the disk (see Fig. 1). The remaining stars are considered halo stars. We carefully exclude all substructures, including streams and overdensities, using the friend-of-friend method (Sun2025ApJ...979..213S; Wang2022MNRAS.513.1958W; Yang2019ApJ...880...65Y). We further remove stars with velocity errors >100>100 km/s and perform $3\sigma$ clipping to remove the outliers. We preserved a refined halo sample comprising 14,497 K-giants, with 11,415 in the northern hemisphere and 3,082 in the southern hemisphere. The sample spans a vast area in the northern hemisphere, up to $r\lesssim 50$ kpc.
@@ 2.2 Correction of the selection function
We corrected for the selection effects of the LAMOST survey following previous studies (Liu2017RAA....17...96L), where the photometric data are assumed to be complete within their limiting magnitude. In the colour-magnitude plane $(c,m)$ along each line of sight $(l,b)$, the photometric density $\nu_{\rm ph}$ is equal to the spectroscopic density $\nu_{\rm sp}$ times the inverse of the selection function $S^{-1}$
$\nu_{\rm ph}(d|c,m,l,b)=\nu_{\rm sp}(d|c,m,l,b)S^{-1}(c,m,l,b),$ (1)
The spectroscopic density $\nu_{\rm sp}(d|c,m,l,b)$ is calculated using the kernel density estimation (KDE) method:
@@
The selection function $S^{-1}(c,m,l,b)$ is evaluated from
@@
The stellar density along a given line of sight is thus:
@@
The selection-corrected average density distribution of the stellar halo in the $(R_{\rm gc},z_{\rm gc})$ plane ($\bar{\nu}_{\rm ph}(R_{\rm gc},z_{\rm gc})$) can then be constructed (Xu2018MNRAS.473.1244X; Yang2022AJ....164..241Y).
We binned the stars observed spectroscopically in the $(R_{\rm gc},z_{\rm gc})$ plane and obtained the number of stars in each bin $n_{\rm sp}(R_{\rm gc},z_{\rm gc})$. For any stars in the bin $i$, we calculate their photometric weight:
@@
where $V=2\pi R_{i}dRdz$ is the volume of the bin. We consider that all the stars in a $1.0\times 1.0$ kpc bin have the same weight.
We present our sample of k-giants in the northern hemisphere in Fig.2. The observations close to the disk plane and at $r\gtrsim 30$ kpc are less complete; thus, the stars have relatively higher weights. With weights correcting for their bias from observational selection, the sample in the northern hemisphere can be considered representative of the halo stars in 6D position-velocity distributions.
@@ 2.3 Correction of the LMC bias
The influence of LMC on the stellar halo at $r<50$ kpc is small. We bin the data in the $R_{\rm gc}-z_{\rm gc}$ plane with a bin size of $10\times 10$ kpc and evaluate their mean velocities $v_{x}$, $v_{y}$, $v_{z}$ within each bin. We observe a minor net velocity in $v_{z}$ and $v_{x}$, around $\sim 10-20$ km/s, whereas a system in perfect dynamical equilibrium should have these velocities at 0. These offsets are presumably caused by the LMC’s influence (Erkal2020MNRAS.498.5574E). We adjusted for this by deducting the mean $v_{z}$ and $v_{x}$ velocities from each star’s velocity, ensuring zero mean velocities in every bin. Although the LMC’s effect is limited in the areas under examination, correcting for it does not significantly alter our final results in constraining the 3D distribution of the DM halo. However, the agreement between the velocity distributions of the model and the data improves slightly when this correction is applied, as expected.
@@ 3 Model construction
Our sample serves as a good approximation of being stationary; i.e., the stars’ DF $f(\boldsymbol{x},\boldsymbol{v})$ will statistically not change when orbiting them in the correct gravitational potential. Driven by data insights, we presented a novel dynamical model (Zhu2025A&A...703A..43Z) aimed at constraining the gravitational potential by directly comparing the observed data distribution $f_{\rm data}(\boldsymbol{x},\boldsymbol{v})$ with the distribution derived from a model superposed by the orbits of these stars integrated within the potential $f_{\rm model}(\boldsymbol{x},\boldsymbol{v}|\rm potential)$. This approach, named the empirical triaxial orbit superposition method - based on minimal assumptions of dynamical models that the system is stationary - overcomes key limitations of previous models. The method has been validated and proves effective at uncovering the 3D DM distribution in galaxies similar to the Milky Way. We have applied it to mock data created from Auriga simulations, selecting three galaxies with different 3D shapes of the underlying DM halo. For all three cases, our model accurately recovered the 3D DM distribution, including the axis ratios and the radial density profile. For one case with a twisted DM halo from the inner to outer regions, our model succeeded in capturing the radial variation of DM halo shape.
@@ 3.1 Empirical triaxial orbit superposition method
Here we briefly introduce the main steps to create the empirical orbit superposition model: (1) we build a model for the gravitational potential with a few free parameters; (2) given one set of parameters, we use the 6D phase space information of stars in the northern hemisphere as initial conditions, calculate their stellar orbits in the gravitational potential; (3) we superpose the stellar orbits with their weights given by the selection function correction and construct stellar density distributions and velocity distributions from the orbit superposition model; (4) we compare the data and model to calculate the likelihood or $\chi^{2}$ of each model; (5) we explore the parameter grid of the gravitational potential and find the best-fitting models with the maximum likelihood or least $\chi^{2}$.
We construct a model of the gravitational potential by including a barred bulge, a thin disk, a thick disk, and a triaxial DM halo with the following density profiles.
@@
$\rho_{\rm disk}=\frac{\Sigma_{0}}{2h}\times\exp(-[\frac{R}{R_{d}}]^{1/n}-\frac{R_{\rm cut}}{R})\times\exp(-\frac{|z|}{h});$ (7)
We fix the density distribution of the bulge and disks obtained from previous studies (Bland-Hawthorn2016): with $M_{*,\rm bulge}=1.6\times 10^{10}M_{\odot}$, scale radius $a=3.5$ kpc, axis ratios $p=0.44$, $q=0.31$ for the bulge, $M_{*,\rm thin}=3.16\times 10^{10}M_{\odot}$, scale radius $R_{d}=2.6$ kpc, scale height $h=0.3$ kpc, inner cutoff radius $R_{\rm cut}=7$ kpc (Lian2024NatAs...8.1302L), Sersic index $n=1$ for the thin disk and $M_{*,\rm thick}=6\times 10^{9}M_{\odot}$, scale radius $R_{d}=2.0$ kpc, scale height $h=0.9$ kpc, Sersic index $n=1$ for the thick disk  [^1] We have tried alternative models, including a model combining a disk without an inner cutoff and a Sersic bulge, as well as a model with circular velocity at the solar position ($r=8.2$ kpc (Bland-Hawthorn2016)) fixed at $V_{c}=235$ km/s (McMillan2017MNRAS.465...76M); neither of them makes noticeable differences in our final results on the DM shape. .
We adopt a flexible triaxial generalised NFW model for the DM halo, allowing free orientations of the DM halo and a variable 3D shape as a function of radius, generally following (Vasiliev2021MNRAS.501.2279V):
@@
in which $\tilde{r}=\sqrt{X^{2}+(Y/p_{\rm DM})^{2}+(Z/q_{\rm DM})^{2}}$, $p_{\rm DM}$, and $q_{\rm DM}$ are the ratios of the axes. The coordinates $\boldsymbol{X}\equiv\{X,Y,Z\}$ are related to the usual Galactocentric Cartesian coordinates $\boldsymbol{x}_{\rm gc}\equiv\{x_{\rm gc},y_{\rm gc},z_{\rm gc}\}$ by a rotation matrix $\boldsymbol{X}=\mathsf{R}\boldsymbol{x}_{\rm gc}$ parameterized by three Euler angles $\alpha_{q}$, $\beta_{q}$, $\gamma_{q}$. The $Z$-axis is tilted by the angle $\beta_{q}$ relative to the $z_{\rm gc}$-axis. Meanwhile, the angles $\alpha_{q}$ ($\gamma_{q}$) define the rotation between the $x_{\rm gc}(X)$ axis and the intersection line of the $x_{\rm gc}y_{\rm gc}$ and $XY$ planes.
The tilt angle $\beta_{q}$ highly degenerates with the vertical axis ratio $q_{\rm DM}$, we fix $\beta_{q}=0^{\circ}$ and allow $q_{\rm DM}>1$ in our model, the $Z$ axis is thus always defined as perpendicular to the stellar disk. Because the direct observations are not complete and we have averaged the information in the azimuthal direction, our model cannot constrain the azimuthal orientation of the DM halo. We fix $\alpha_{q}=0^{\circ}$ in our fiducial model.
For the radial profile, we fixed $\alpha=1$, $\beta=3$, chose the outer cutoff radius $r_{\rm cut}=500$ kpc and the cutoff strength $\xi=5$. We were left with five free parameters in the DM halo: $\rho_{0}$, $r_{s}$, $\gamma$, $p_{\rm DM}$, and $q_{\rm DM}$.
When needed, we allow $p_{\rm DM}$ and $q_{\rm DM}$ to vary as a function of radius, with:
@@
$q_{\rm DM}(r<10)=q_{\rm in}$ (12)
where $r=\sqrt{X^{2}+Y^{2}+Z^{2}}$, $p_{\rm in}$, $p_{\rm out}$, $q_{\rm in}$, $q_{\rm out}$, and the scale radius $r_{q}$ are allowed to be free parameters.
Given each set of parameters in the gravitational potential, we integrate the stellar orbits using the publicly released AGAMA package (Vasiliev2019MNRAS.482.1525V) [^2] https://github.com/GalacticDynamics-Oxford/Agama for the 11415 stars in the northern hemisphere. We integrate 10 orbital periods for each star and withdraw 1000 particles from each orbit with equal time steps, and each particle is given the weight of the star that initialises the orbit. We superpose particles drawn from the orbits together, thus obtaining an orbit superposed model, which numerically represents the DF in 6D phase space of the Milky Way stellar halo.

### v4

@@ 5 Department of Astronomy, Westlake University, Hangzhou, Zhejiang 310030, China
A vertically orientated dark matter halo marks a flip of the Galactic disc
@@
Unveiling the 3D shape of the Milky Way’s dark-matter halo is critical to understanding its formation history. We created an innovative dynamical model that makes minimal assumptions about the internal dynamical structures and accommodates a highly flexible triaxial DM halo. By applying the method to 6D phase-space data of K-giant stars from LAMOST + Gaia, we robustly determined the 3D dark-matter distribution of the Milky Way out to approximately $50$ kpc. We discovered a triaxial, nearly oblate dark-matter halo with $q_{\rm DM}=Z/X=0.92\pm 0.08$, $p_{\rm DM}=Y/X=0.8\pm 0.2$ on average within 50 kpc, where the $Z$ axis is defined perpendicular to the stellar disc. The axes ratio $q_{\rm DM}>p_{\rm DM}$ is strongly preferred; the long-intermediate axis plane of the dark-matter halo is unexpectedly vertical to the Galactic disc, yet aligned with the ‘plane of satellites’. This striking configuration suggests that the Galactic disc (and the inner halo) has flipped, likely torqued by minor mergers, from an original alignment with the outer dark-matter halo and satellite plane, as is supported by Milky Way analogues from Auriga and TNG50. By allowing $q_{\rm DM}(r)$ and $p_{\rm DM}(r)$ to vary with radii, we find tentative evidence that the dark-matter halo is twisted. This agrees alignment with the disc in the inner regions and transitions to a vertical orientation at $r\gtrsim 20$ kpc, supporting the disc flip scenario prediction. Such disc reorientation is non-trivial, yet its physical mechanism is straightforward to comprehend and naturally originates a vertical satellite plane. Our findings offer a unified framework that links dark-matter halo orientation, satellite alignment, and disc evolution, reinforcing the internal consistency of the Milky Way in the $\Lambda$ cold dark matter model.
@@ 1 Introduction
Determining the exact shape of the DM halo from observations, even for the Milky Way, is a challenge. One major method of probing the DM distribution is to model the motions of tracers, which reflect the underlying gravitational potential. Stellar streams, viewed as simple orbital pathways that trace the halo, played a significant role in constraining the shape of the Milky Way DM halo (Bovy et al., 2016; Bowden et al., 2015; Koposov et al., 2010; Palau and Miralda-Escudé, 2023; Law and Majewski, 2010). However, the most commonly used dynamically cold streams, for example Pal 5 and GD–1, trace only the inner halo at $r\lesssim 20$ kpc. And a single stream is easily disturbed (Nibauer et al., 2024) and may lead to biassed results (Vasiliev et al., 2021).
@@
Halo stars – with the full 6D phase-space information observed – span the entirety of space and, statistically, should exhibit strong power in constraining the underlying 3D DM distribution. However, these stars exist with complex orbital dynamics, necessitating accurate dynamical modelling. At present, the Milky Way halo is analysed using two primary types of dynamical models: the Jeans model and the distribution function (DF)-based dynamical model. These models yield divergent results regarding the shape of the DM halo, ranging from oblate with $q\sim 0.4-0.7$ (Loebman et al., 2014; Hattori et al., 2021; Li and Binney, 2022), to nearly spherical with $q\sim 0.9-1$ (Wegg et al., 2019; Zhang et al., 2025), and to prolate with $q\sim 1.3$ (Posti and Helmi, 2019). All these models are restricted by axisymmetric assumptions and incorporate ad hoc assumptions about the intrinsic dynamical structures. Despite extensive research over the years, the true shape of the Milky Way DM halo remains debated (Hunt and Vasiliev, 2025).
@@
Due to the large uncertainty surrounding the DM halo shape of the Milky Way from observations, its link to the spatial arrangement of satellite galaxies has not been seriously considered. The satellite galaxies in the Milky Way are not distributed isotropically but rather form a thin plane vertical to the Galactic disc (Lynden-Bell, 1976; Kroupa et al., 2005; Pawlowski, 2018; Sawala et al., 2023). The thickness of this plane of satellites has been extensively studied; it was previously thought to be rare but has now become more frequently observed through cosmological simulations (Sawala et al., 2023). The cause of its perpendicular alignment to the Galactic disc remains uncertain (Vasiliev et al., 2021). Typically, both the stellar disc and satellite system align parallel to the DM halo, sharing a similar net angular momentum direction, leading to a co-alignment of the stellar disc and satellite plane (Shao et al., 2016). However, in systems resembling the Milky Way, in which the satellite system lies perpendicular to the stellar disc, the DM halo might exhibit a twist, being co-aligned with the stellar disc in the inner regions but aligned vertically in the outer regions (Shao et al., 2021). Precisely measuring the DM shape as a function of radius could provide essential insights into the origin of the vertical satellite plane and elucidate its relationship with the large-scale matter distribution. Additionally, this suggests that the DM halo of the Milky Way may be triaxial, with a shape that changes from the inner to outer regions, necessitating dynamic models with more flexible DM halos.
@@
Recently, a triaxial Schwarzwald model that numerically represents the DF of tracers by superposition of stellar orbits has been applied to the Milky Way halo, indicating a triaxial, nearly prolate halo with a large tilt angle (Dillamore and Sanders, 2025). Nevertheless, there exists a strong degeneracy between the axis ratios of the DM halo and the tilt angle. This model was mainly constrained by the shape of a tilted stellar halo (Han et al., 2022a) associated with the Gaia-Sausage-Enceladus (GSE) (Helmi et al., 2018; Belokurov et al., 2018).
@@
In a previous paper (Zhu et al., 2025), we presented and validated an empirical triaxial orbit-superposition model tailored for the Milky Way halo observed by the Large Sky Area Multi-Object Fibre Spectroscopic Telescope (LAMOST) and the Global Astrometric Interferometer for Astrophysics (Gaia). This approach accommodates a highly flexible triaxial DM halo, where the axis ratios are allowed to vary with radius, and it uses the full 6D phase-space information from the observations. Compared to traditional made-to-measure and Schwarzschild methods, this method does not theoretically sample the orbit libraries or determine the orbit weights by fitting the model to data; instead, the orbit library and orbit weights are entirely determined from the observations. The results are thus highly data-driven.
@@
In this paper, we apply the empirical triaxial orbit-superposition model to the Milky Way halo, utilising 6D phase-space data from LAMOST and Gaia, and robustly determine the 3D DM distribution of the Milky Way, extending to approximately 50 kpc. We then compare the findings with Milky Way analogues from the Auriga and TNG50 cosmological simulations and establish a coherent scenario linking the DM halo shape and satellite arrangement. The paper is organised in the following way: we describe the details of observational data in Section 2, introduce model construction in Section 3, and show the results in Section 4. We compare the 3D shape of the DM halo and the arrangement of the Milky Way’s satellite system with cosmological simulations in Section 5. We discuss in Section 6 and conclude in Section 7.
@@ 2.1 Data preparation
Figure 1: Density of stars in the parameter phase of circularity, $\lambda_{z}$, versus metallicity, [Fe/H]. We first removed the disc stars with a circularity of $\lambda_{z}>0.8$, then also removed the group of stars with $\lambda_{z}>0.3$ and $\rm Fe/H]>-1.2$. The rest were kept as halo stars in our sample.
Figure 2: K giants in our final smooth halo sample in the northern hemisphere, coloured by the weights of particles obtained by correction of the selection function. The sample can be taken as complete in the $R_{\rm gc}-z_{\rm gc}$ space after correction of the selection function.
We obtained the 3D positions and 3D velocities of halo stars by combining observations from LAMOST and Gaia. LAMOST has gathered spectroscopic data for approximately 10 million stars within the Milky Way. We used k giants from LAMOST DR9; the distance is accurately measured with a relative uncertainty of $\sim 15\%$ (Zhang et al., 2023). By cross-matching with Gaia DR3, we obtained a sample of 619,284 k-giants with 3D velocity measurements; the typical errors in the radial velocity, $v_{r}$, tangential velocity, $v_{\phi}$, and $v_{\theta}$ are 10, 30, and 30 km/s, respectively. We excluded disc stars chemodynamically to avoid biassing the spatial distribution of the halo. We first removed all of the stars in the cold disc with $\lambda_{z}>0.8$, then removed the group of stars with $\lambda_{z}>0.3$ and $\rm Fe/H]>-1.2$ that are mostly within 20 kpc, not in dynamical equilibrium, and likely to be induced from the disc (see Fig. 1). The remaining stars were considered halo stars. We carefully excluded all substructures, including streams and overdensities, using the friend-of-friend method (Sun et al., 2025; Wang et al., 2022; Yang et al., 2019). We further removed stars with velocity errors of >100>100 km/s and performed $3\sigma$ clipping to remove the outliers. We preserved a refined halo sample comprising 14,497 K-giants, with 11,415 in the northern hemisphere and 3,082 in the southern hemisphere. The sample spans a vast area in the northern hemisphere, up to $r\lesssim 50$ kpc.
@@ 2.2 Correction of the selection function
We corrected for the selection effects of the LAMOST survey following previous studies (Liu et al., 2017), in which the photometric data were assumed to be complete within their limiting magnitude. In the colour-magnitude plane $(c,m)$ along each line of sight $(l,b)$, the photometric density, $\nu_{\rm ph}$, is equal to the spectroscopic density, $\nu_{\rm sp}$, times the inverse of the selection function, $S^{-1}$,
$\nu_{\rm ph}(d|c,m,l,b)=\nu_{\rm sp}(d|c,m,l,b)S^{-1}(c,m,l,b).$ (1)
The spectroscopic density, $\nu_{\rm sp}(d|c,m,l,b)$, was calculated using the kernel density estimation (KDE) method:
@@
The selection function, $S^{-1}(c,m,l,b)$, was evaluated from
@@
The stellar density along a given line of sight is thus
@@
The selection-corrected average density distribution of the stellar halo in the $(R_{\rm gc},z_{\rm gc})$ plane ($\bar{\nu}_{\rm ph}(R_{\rm gc},z_{\rm gc})$) could then be constructed (Xu et al., 2018; Yang et al., 2022).
We binned the stars observed spectroscopically in the $(R_{\rm gc},z_{\rm gc})$ plane and obtained the number of stars in each bin $n_{\rm sp}(R_{\rm gc},z_{\rm gc})$. For each of the stars in the bin, $i$, we calculated their photometric weight:
@@
where $V=2\pi R_{i}dRdz$ is the volume of the bin. We considered all the stars in a $1.0\times 1.0$ kpc bin to have the same weight.
We present our sample of k-giants in the northern hemisphere in Fig.2. The observations close to the disc plane and at $r\gtrsim 30$ kpc are less complete; thus, the stars have higher weights. With weights correcting for their bias from observational selection, the sample in the northern hemisphere can be considered representative of the halo stars in 6D position-velocity distributions.
@@ 2.3 Correction of the LMC bias
The influence of LMC on the stellar halo at $r<50$ kpc is small. We binned the data in the $R_{\rm gc}-z_{\rm gc}$ plane with a bin size of $10\times 10$ kpc and evaluated their mean velocities, $v_{x}$, $v_{y}$, and $v_{z}$, within each bin. We observe a minor net velocity in $v_{z}$ and $v_{x}$, around $\sim 10-20$ km/s, whereas a system in perfect dynamical equilibrium should have these velocities at 0. These offsets are presumably caused by the LMC’s influence (Erkal et al., 2020). We adjusted for this by deducting the mean $v_{z}$ and $v_{x}$ velocities from each star’s velocity, ensuring zero mean velocities in every bin. Although the LMC’s effect is limited in the areas under examination, correcting for it does not significantly alter our final results in constraining the 3D distribution of the DM halo. However, the agreement between the velocity distributions of the model and the data improves slightly when this correction is applied, as was expected.
@@ 3 Model construction
Our sample serves as a good approximation of being stationary, i.e. the stars’ DF $f(\boldsymbol{x},\boldsymbol{v})$ will statistically not change when orbiting them in the correct gravitational potential. Driven by data insights, we presented a novel dynamical model (Zhu et al., 2025) aimed at constraining the gravitational potential by directly comparing the observed data distribution, $f_{\rm data}(\boldsymbol{x},\boldsymbol{v})$, with the distribution derived from a model superposed by the orbits of these stars integrated within the potential $f_{\rm model}(\boldsymbol{x},\boldsymbol{v}|\rm potential)$. This approach, named the empirical triaxial orbit superposition method – based on minimal assumptions of dynamical models that the system is stationary – overcomes key limitations of previous models. The method has been validated and proves effective at uncovering the 3D DM distribution in galaxies similar to the Milky Way. We have applied it to mock data created from Auriga simulations, selecting three galaxies with different 3D shapes of the underlying DM halo. In all three cases, our model accurately recovers the 3D DM distribution, including the axis ratios and the radial density profile. For one case with a twisted DM halo from the inner to outer regions, our model succeeds in capturing the radial variation in the DM halo shape.
@@ 3.1 Empirical triaxial orbit superposition method
Here we briefly introduce the main steps to create the empirical orbit superposition model: (1) we built a model for the gravitational potential with a few free parameters; (2) given one set of parameters, we used the 6D phase space information of stars in the northern hemisphere as initial conditions, and calculated their stellar orbits in the gravitational potential; (3) we superposed the stellar orbits with their weights given by the selection function correction and constructed stellar density distributions and velocity distributions from the orbit superposition model; (4) we compared the data and model to calculate the likelihood or $\chi^{2}$ of each model; (5) we explored the parameter grid of the gravitational potential and found the best-fitting models with the maximum likelihood or least $\chi^{2}$.
We constructed a model of the gravitational potential by including a barred bulge, a thin disc, a thick disc, and a triaxial DM halo with the following density profiles:
@@
$\rho_{\rm disc}=\frac{\Sigma_{0}}{2h}\times\exp(-[\frac{R}{R_{d}}]^{1/n}-\frac{R_{\rm cut}}{R})\times\exp(-\frac{|z|}{h}).$ (7)
We fixed the density distribution of the bulge and discs obtained from previous studies (Bland-Hawthorn and Gerhard, 2016): with $M_{*,\rm bulge}=1.6\times 10^{10}M_{\odot}$, a scale radius of $a=3.5$ kpc, axis ratios of $p=0.44$, $q=0.31$ for the bulge, $M_{*,\rm thin}=3.16\times 10^{10}M_{\odot}$, a scale radius of $R_{d}=2.6$ kpc, a scale height of $h=0.3$ kpc, an inner cut-off radius of $R_{\rm cut}=7$ kpc (Lian et al., 2024), a Sersic index of $n=1$ for the thin disc and $M_{*,\rm thick}=6\times 10^{9}M_{\odot}$, a scale radius of $R_{d}=2.0$ kpc, a scale height of $h=0.9$ kpc, and a Sersic index of $n=1$ for the thick disc [^1] We tried alternative models, including a model combining a disc without an inner cut-off and a Sersic bulge, as well as a model with circular velocity at the solar position ($r=8.2$ kpc (Bland-Hawthorn and Gerhard, 2016)) fixed at $V_{c}=235$ km/s (McMillan, 2017); none of them makes noticeable differences to our final results for the DM shape. .
We adopted a flexible triaxial generalised NFW model for the DM halo, allowing free orientations of the DM halo and a variable 3D shape as a function of radius, generally following (Vasiliev et al., 2021):
@@
in which $\tilde{r}=\sqrt{X^{2}+(Y/p_{\rm DM})^{2}+(Z/q_{\rm DM})^{2}}$, $p_{\rm DM}$, and $q_{\rm DM}$ are the ratios of the axes. The co-ordinates $\boldsymbol{X}\equiv\{X,Y,Z\}$ are related to the usual Galactocentric Cartesian co-ordinates $\boldsymbol{x}_{\rm gc}\equiv\{x_{\rm gc},y_{\rm gc},z_{\rm gc}\}$ by a rotation matrix, $\boldsymbol{X}=\mathsf{R}\boldsymbol{x}_{\rm gc}$, parameterised by three Euler angles, $\alpha_{q}$, $\beta_{q}$, and $\gamma_{q}$. The $Z$ axis is tilted by the angle $\beta_{q}$ relative to the $z_{\rm gc}$ axis. Meanwhile, the angles $\alpha_{q}$ ($\gamma_{q}$) define the rotation between the $x_{\rm gc}(X)$ axis and the intersection line of the $x_{\rm gc}y_{\rm gc}$ and $XY$ planes.
The tilt angle $\beta_{q}$ highly degenerates with the vertical axis ratio, $q_{\rm DM}$. We fixed $\beta_{q}=0^{\circ}$ and allowed $q_{\rm DM}>1$ in our model. The $Z$ axis is thus always defined as perpendicular to the stellar disc. Because the direct observations are not complete and we have averaged the information in the azimuthal direction, our model cannot constrain the azimuthal orientation of the DM halo. We fixed $\alpha_{q}=0^{\circ}$ in our fiducial model. In this model with fixed $\beta_{q}$ and $\alpha_{q}$, the axis ratio $q_{\rm DM}$ and $p_{\rm DM}$ will reflect the DM halo orientation.
For the radial profile, we fixed $\alpha=1$, $\beta=3$, and chose the outer cut-off radius $r_{\rm cut}=500$ kpc and the cut-off strength $\xi=5$. We were left with five free parameters in the DM halo: $\rho_{0}$, $r_{s}$, $\gamma$, $p_{\rm DM}$, and $q_{\rm DM}$.
When needed, we allowed $p_{\rm DM}$ and $q_{\rm DM}$ to vary as a function of radius, with
@@
$q_{\rm DM}(r<10)=q_{\rm in},$ (12)
where $r=\sqrt{X^{2}+Y^{2}+Z^{2}}$, $p_{\rm in}$, $p_{\rm out}$, $q_{\rm in}$, $q_{\rm out}$, and the scale radius, $r_{q}$, were allowed to be free parameters.
Given each set of parameters in the gravitational potential, we integrated the stellar orbits using the publicly released AGAMA package (Vasiliev, 2019) [^2] https://github.com/GalacticDynamics-Oxford/Agama for the 11415 stars in the northern hemisphere. We integrated ten orbital periods for each star and withdrew 1000 particles from each orbit with equal time steps, and each particle was given the weight of the star that initialises the orbit. We superposed particles drawn from the orbits together, obtaining an orbit superposed model that numerically represents the DF in 6D phase space of the Milky Way stellar halo.